# Exact Mobility Edges in a Disorder-Free Dimerized Stark Lattice with Effective Unbounded Hopping

### v2

@@ F.2 Input-Output Formalism and Transmission Spectrum
To connect the theoretical Hamiltonian to the experimental transmission signal, we employ the Heisenberg-Langevin formalism [Walls_2008]. The time evolution of the annihilation operators for the cavity modes, collected in the vector $\mathbf{\Psi}=(\dots,a_{n},\dots,b_{n},\dots)^{T}$, is given in the Heisenberg picture by
@@
$\tilde{a}_{n}=a_{n}e^{i(\omega_{A}+n\omega_{\text{mod}})t},\quad\tilde{b}_{n}=b_{n}e^{i(\omega_{A}+n\omega_{\text{mod}})t}.$ (48)
@@
where $H_{\text{eff}}$ is the static matrix representation of the effective Hamiltonian derived in Sec. F.1.
@@
$s_{\text{in}}^{\text{rot}}(t)=s_{\text{in}}^{\text{lab}}(t)e^{i\omega_{A}t}=\mathcal{E}e^{-i(\omega_{\text{laser}}-\omega_{A})t}.$ (50)
By locking the laser frequency to the central resonance of resonator A ($\omega_{\text{laser}}=\omega_{A}$), the time dependence vanishes, yielding a constant input vector $\tilde{\mathbf{S}}_{\text{in}}$ with a single nonzero entry corresponding to the $\tilde{a}_{0}$ mode.
@@
The steady-state output field is determined by the input-output boundary condition $\tilde{\mathbf{S}}_{\text{out}}=\tilde{\mathbf{S}}_{\text{in}}-\sqrt{\kappa_{\text{ex}}}\mathbf{\Phi}_{\text{ss}}$ (through port) or $\tilde{\mathbf{S}}_{\text{out}}=\sqrt{\kappa_{\text{ex}}}\mathbf{\Phi}_{\text{ss}}$ (drop port). The optical spectrum analyzer (OSA) measures the power at each frequency mode $\omega_{n}$, which is proportional to the mode occupation $|\tilde{a}_{n}|^{2}$. Since $|\tilde{a}_{n}|^{2}=|a_{n}|^{2}$, the rotating frame photon number directly corresponds to the laboratory frame observable.

### v3

@@ F.2 Input-output formalism and transmission spectrum
To connect the theoretical Hamiltonian to the experimental transmission signal, we employ the Heisenberg-Langevin formalism [70]. The time evolution of the annihilation operators for the cavity modes, collected in the vector $\mathbf{\Psi}=(\dots,a_{n},\dots,b_{n},\dots)^{T}$, is given in the Heisenberg picture by
@@
$\tilde{a}_{n}=a_{n}e^{i(\omega_{\text{A}}+n\omega_{\text{mod}})t},\quad\tilde{b}_{n}=b_{n}e^{i(\omega_{\text{A}}+n\omega_{\text{mod}})t}.$ (48)
@@
where $H_{\text{eff}}$ is the static matrix representation of the effective Hamiltonian derived in Appendix F.1.
@@
$s_{\text{in}}^{\text{rot}}(t)=s_{\text{in}}^{\text{lab}}(t)e^{i\omega_{\text{A}}t}=\mathcal{E}e^{-i(\omega_{\text{laser}}-\omega_{\text{A}})t}.$ (50)
By locking the laser frequency to the central resonance of resonator A ($\omega_{\text{laser}}=\omega_{\text{A}}$), the time dependence vanishes, yielding a constant input vector $\tilde{\mathbf{S}}_{\text{in}}$ with a single nonzero entry corresponding to the $\tilde{a}_{0}$ mode.
@@
The steady-state output field is determined by the input-output boundary condition $\tilde{\mathbf{S}}_{\text{out}}=\tilde{\mathbf{S}}_{\text{in}}-\sqrt{\kappa_{\text{ex}}}\mathbf{\Phi}_{\text{ss}}$ (through port) or $\tilde{\mathbf{S}}_{\text{out}}=\sqrt{\kappa_{\text{ex}}}\mathbf{\Phi}_{\text{ss}}$ (drop port). The optical spectrum analyzer measures the power at each frequency mode $\omega_{n}$, which is proportional to the mode occupation $|\tilde{a}_{n}|^{2}$. Since $|\tilde{a}_{n}|^{2}=|a_{n}|^{2}$, the rotating frame photon number directly corresponds to the laboratory frame observable.